# ClinKD: Cross-Modal Clinical Knowledge Distiller For Multi-Task Medical Images

### v3

@@ 4.4.3 Comparison with Other Methods
Comparison with BiRD + RoPE-Mixed [12]. The RoPE-Mixed uses frequencies for both axes as learnable network parameters, effectively handling image features in diagonal direction. It achieves the best performance in different visual scenarios [48, 49, 23]. Hence we compare our MCG-RoPE with RoPE-Mixed. As shown in Table 3, by adding RoPE-Mixed to BiRD [14], great achievements take place on all modalities (CT, MR, X-ray, etc.). Compared with RoPE-Mixed [12], our MCG-RoPE outperforms it by 4.95%, 2.94%, 5.78% and 2.60% on VG, ROC, RC and MIA tasks, respectively. For more detailed performance on the 8 modalities, MCG-RoPE still achieves the best improvements most of the time.
Comparison with BiRD + $V_{k}D$ [30]. $V_{k}D$ is a novel method that can improve knowledge distillation. This work utilizes a projection layer for maximizing the knowledge transfer to the student backbone. Many experiments [8, 33, 31, 13] have been made to prove the best performance of this method. Compared with BiRD with $V_{k}D$ in the Table 3, Pseudo-KD shows its advantages on multi-task fine-grained Med-VQA scenarios, outperforming $V_{k}D$ almost on all modalities.
Figure 5: Visualization of attentions on GC and VG tasks. The ClinKD generates better answers compared with BiRD [14].
@@ 4.4.4 Case Study
Case study on Med-GRIT-Test30 [39, 7]. Figure 5 illustrates the visualization of the attention map and a comparison among BiRD [14], ClinKD (ours), and the ground truth. The grounding of ClinKD is more accurate because MCG-RoPE can improve image-text alignment, and at the same time, Pseudo-KD enables ClinKD to better adapt to medical knowledge.
Case study on Additional VQA-RAD [16]. As shown in Figure 6, we conducted an experiment on VQA-RAD [16]. Despite the relatively simple annotations, ClinKD provides fine-grained answers with three accurate groundings, demonstrating its precision in response.
To assess the models’ attention to pathological features, we conducted a comparative analysis of three RoPE variants. The results indicate that 2D-RoPE [32] tends to show low attention on pathological regions. In contrast, RoPE-Mixed [12] demonstrates a greater emphasis on diagonal features. Notably, our Med-CLIP Guided RoPE specifically targets and enhances the model’s attention to pathological regions, thereby facilitating more accurate alignment between text and images, especially in the identification of relevant medical features.
Figure 6: An example from VQA-RAD [16] and the performance of RoPE variants.
4.4.5 Limitation
Although our method alleviates, to some extent, the issue of insufficient prior knowledge and achieves impressive state-of-the-art performance on the Med-GRIT-270k [39, 7] dataset, the model still exhibits hallucinations. Furthermore, through comparative evaluations on the LLaVA-Med-qa0.2k [46] dataset, our approach remains insufficient to enable the model to emerge with strong generalization capabilities in Med-VQA domain. Creating high-quality Med-VQA datasets and training medical prior knowledge before freezing the visual encoder could significantly enhance the model’s generalization ability on medical data.
@@ 5 Conclusion
In this work, we propose ClinKD, a novel medical distiller centered on Pseudo-Labels Medical Knowledge Distillation and Reflective Correction Training. In addition, we propose the Med-CLIP Guided RoPE to improve image-text alignment. They are designed to enhance the performance on fine-grained multi-task datasets in Med-VQA domain. Extensive experiments demonstrate that our Med-CLIP Guided RoPE achieves superior image-text alignment, while the Pseudo-Labels Medical Knowledge Distillation effectively bridges prior medical knowledge gaps. These mechanisms synergistically enhance the model’s medical knowledge adaptation capabilities.
@@
6 Acknowledgments
This work was supported by the Natural Science Foundation of Guangdong Province (No. 2023A1515010673), in part by the Shenzhen Science and Technology Innovation Bureau key project (No. JSGG20220831110400001, No. CJGJZD20230724093303007,KJZD20240903101259001), in part by Shenzhen Medical Research Fund (No. D2404001), in part by Shenzhen Engineering Labora- tory for Diagnosis & Treatment Key Technologies of Interventional Surgical Robots (XMHT20220104009), and the Key Laboratory of Biomedical Imaging Science and System, CAS, for the Research platform support.
Supplementary Material
A1 Datasets in the Experiments
A1.1 Med-GRIT-270k [14]
Instance-level Meta Information. Initially, approximately 60,000 biomedical image-mask pairs were randomly sampled from the SA-Med2D-20M dataset  [39, 7], considering modality diversity and redundancy. Notably, the original dataset contains numerous 2D slices extracted from 3D volumes, which could lead to high similarity and redundancy among images. Next, instance locations were determined based on instance-level masks. Specifically, spatial coordinates of each instance were represented in textual form as $[X_{topleft},Y_{topleft},X_{bottomright},Y_{bottomright}]$, and subsequently normalized to the range [0, 1]. Finally, each image was annotated with additional meta-information, including imaging modality, scanned anatomical region, orientation, and instance coordinates.
Obtaining Image Captions. Carefully designed prompts, combined with the meta-information, were provided to ChatGPT to extract global context information for each image.
Details of Instruction-Tuning Data. Spatial understanding is manifested through various task formats. This primarily encompasses two distinct types and their corresponding task names: (i) Region-in and Text-out: Referring Object Classification (ROC), Referring Captioning (RC), (ii) Text-in and Region-out: Visual Grounding (VG), and (iii) Text-in and Text-out: Medical Image Analysis (MIA). To reduce ambiguity and enhance the model’s capability for fine-grained visual comprehension, some essential strategies are adopted. The special tokens (<ref >and </ref>) are introduced, marking the content referred to by the bounding box. This aptly associates bounding boxes with their corresponding descriptive words or sentences. Subsequently, we instructed ChatGPT to design a question and answer for each task. Finally, We mapped the coordinates within the range [0, 1000] and reformatted them as $(X_{topleft},Y_{topleft})$, $(X_{bottomright},Y_{bottomright})$. To differentiate between detection strings and regular text strings, two special tokens (<box>and </box>) are appended at the start and end of the bounding box string, respectively.
The proportions of various modalities are shown in the Figure A1 and the Figure A2 shows an example of instruction-tuning conversations.
Figure A1: The proportions of various modalities in the Med-GRIT-270k [14]
A1.2 LLaVA-Med-qa0.2k [18]
The LLaVA-Med-qa0.2k is primarily used for evaluating LLaVA-Med [18]. It is an evaluation dataset consisting of 193 novel questions was constructed. Specifically, 50 previously unseen image-caption pairs were randomly selected from the PMC-15M dataset. Conversational questions were produced using the same self-instruct data generation pipeline employed in the second training stage. In contrast, detailed description questions were randomly chosen from a predefined question set, aiming to prompt comprehensive and detailed responses.
Figure A2: The example of instruction-tuning data.
A2 Training Configurations
We use LLaMA-Factory to train Qwen2-VL and here are some information about model configurations and training configurations.
Category Key Value
Formatting ShareGPT
Columns messages messages
images images
Tags role_tag role
content_tag content
user_tag user
assistant_tag assistant
Table A1: Qwen2-VL [34] Data Configuration Summary.
Parameter Value
Model Architecture
Hidden size 3584
Hidden layers 28
Attention heads 28
Key/Value heads 4
Intermediate size 18944
Max position embeddings 32768
Vocab size 152,064
RMS norm epsilon 1e-06
Training/Regularization
Initializer range 0.02
Attention dropout 0.0
Sliding window size 32768
Special Tokens
BOS token ID 151643
EOS token ID 151645
Image token ID 151655
Video token ID 151656
Vision start token ID 151652
Vision end token ID 151653
Vision Configuration
Input channels 3
Spatial patch size 14
Table A2: The Architecture and Configuration Parameters of ClinKD.
Parameter Value
Model Configuration
Base model Qwen2-VL-7B-Instruct
Quantization method bitsandbytes
Fine-tuning type Full
Template qwen2_vl
Training Setup
Batch size 2
Gradient accumulation 8
Max sequence length 2048
Training epochs 3
Learning rate $2\times 10^{-5}$
Warmup steps 0
Optimizer AdamW
Learning schedule Cosine
Memory Management
DeepSpeed stage 3
Offloading Disabled
BF16 computation Enabled
Max grad norm 1.0
Adapter Configuration
LoRA rank 8
LoRA alpha 16
LoRA dropout 0
Trainable layers 2
Advanced Options
BAdam mode Layer (ascending)
BAdam update ratio 0.05
GaLoRe rank 16
SFT stage Supervised Fine-Tuning
Reward modeling Disabled
Table A3: Training Configuration Parameters for ClinKD.
During the knowledge distillation process, additional GPU memory allocation is required to accommodate the teacher model in inference mode. For instance, the Qwen2-VL-7B model consumes approximately 25GB of GPU memory per question-answering iteration. When processing large batches of data, we recommend allocating at least 80GB of GPU memory for the teacher model to ensure smooth operation.

### v4

@@ 4.3.3 Comparison with Other Methods
Comparison with BiRD + RoPE-Mixed heo2024rotarypositionembeddingvision . The RoPE-Mixed uses frequencies for both axes as learnable network parameters, effectively handling image features in diagonal direction. It achieves the best performance in different visual scenarios ade20k ; zhou2018semanticunderstandingscenesade20k ; coco . Hence we compare our MCG-RoPE with RoPE-Mixed. As shown in Table 3, by adding RoPE-Mixed to BiRD huang2024BiRD , great achievements take place on all modalities (CT, MR, X-ray, etc.). Compared with RoPE-Mixed heo2024rotarypositionembeddingvision , our MCG-RoPE outperforms it by 4.95%, 2.94%, 5.78% and 2.60% on VG, ROC, RC and MIA tasks, respectively. For more detailed performance on the 8 modalities, MCG-RoPE still achieves the best improvements most of the time.
Comparison with BiRD + $V_{k}D$ Miles_2024_CVPR . $V_{k}D$ is a novel method that can improve knowledge distillation. This work utilizes a projection layer for maximizing the knowledge transfer to the student backbone. Many experiments DLGAN_CVPR ; spkd ; pkd ; label-kd have been made to prove the best performance of this method. Compared with BiRD with $V_{k}D$ in the Table 3, Pseudo-KD shows its advantages on multi-task fine-grained Med-VQA scenarios, outperforming $V_{k}D$ almost on all modalities.
4.3.4 Few-shot Experiments
Table 4 shows the results of our additional experiments for testing few-shot performance. The evaluation scores shows that our ClinKD achieves the best performance on all datasets. It is can be seen that the BiRD which also utilizes Rotary Position Embedding outperforms LLaVA-Med li2024llava and Med-Flamingo med-fla . This may be caused by the special meaning of 2D-RoPE. For our ClinKD, the modified Med-CLIP Guided RoPE enhances the position information between image tokens and text tokens, enabling better image-text alignment.
Dataset Metric LLaVA-Med li2024llava BiRD huang2024BiRD Med-Flamingo med-fla ClinKD
VQA-RAD lau2018dataset F1 42.91 46.12 41.79 59.21
BLEU-1 43.35 48.23 43.27 57.09
SLAKE liu2021slake F1 44.28 49.25 45.86 58.64
BLEU-1 43.35 48.23 41.98 58.63
Path-VQA he2020pathvqa F1 44.28 49.25 42.56 59.21
Table 4: Comparison with LLaVA-Med li2024llava and BiRD (previous SOTA) huang2024BiRD . The best scores are shown in bold.
@@ 4.3.5 Case Study
Case study on Med-GRIT-Test30 ye2023samed2d20m ; cheng2023sammed2d . Figure 4a illustrates the visualization of the attention map and a comparison among BiRD huang2024BiRD , ClinKD (ours), and the ground truth. The grounding of ClinKD is more accurate because MCG-RoPE can improve image-text alignment, and at the same time, Pseudo-KD enables ClinKD to better adapt to medical knowledge.
Case study on VQA-RAD. As shown in Figure 4b, we conducted an experiment on VQA-RAD lau2018dataset . Despite the relatively simple annotations, ClinKD provides fine-grained answers with three accurate groundings, demonstrating its precision in response.
To assess the models’ attention to pathological features, we conducted a comparative analysis of three RoPE variants. The results indicate that 2D-RoPE su2023roformerenhancedtransformerrotary tends to show low attention on pathological regions. In contrast, RoPE-Mixed heo2024rotarypositionembeddingvision demonstrates a greater emphasis on diagonal features. Notably, our Med-CLIP Guided RoPE specifically targets and enhances the model’s attention to pathological regions, thereby facilitating more accurate alignment between text and images, especially in the identification of relevant medical features.
(a) Case study on Med-GRIT-Test30 ye2023samed2d20m ; cheng2023sammed2d .
(b) Case study on VQA-RAD lau2018dataset .
Figure 4: Case Studies on Med-GRIT-Test30 ye2023samed2d20m ; cheng2023sammed2d and VQA-RAD lau2018dataset .
@@ 5 Conclusion
In this work, we propose ClinKD, a novel medical distiller centered on Adaptive Confidence-Margin Curriculum Pseudo-KD and Reflective Correction Training. In addition, we propose the Med-CLIP Guided RoPE to improve image-text alignment. They are designed to enhance the performance on fine-grained multi-task datasets in Med-VQA domain. Extensive experiments demonstrate that our Med-CLIP Guided RoPE achieves superior image-text alignment, while the Adaptive Confidence-Margin Curriculum Pseudo-KD effectively bridges prior medical knowledge gaps. These mechanisms synergistically enhance the model’s medical knowledge adaptation capabilities.